# Boosting Continual Learning of Vision-Language Models via Mixture-of-Experts Adapters

### v1

@@ 2 Related Works
Figure 2: Overall framework of the proposed method. (a) At the training stage, CLIP’s image and text encoders $(\mathcal{F}_{I},\mathcal{F}_{T})$ take input samples from Task t. In each of transformer blocks, there is a MoE-Adapters, whose input is the tokens $\textbf{{x}}^{t}$ from MHSA. The router takes the task-specific [CLS] token $\textbf{{c}}^{t}$ as input and produces experts’ weights $W_{i}^{t}$ and $W_{i}^{t}$ to combine the expert’s output. DDAS is trained using only images via the MSE loss defined by Eq. 3. (b) At the inference stage, the proposed DDAS determines the data distribution by comparing the distribution $\{d^{t}\}_{t=1}^{T}$ in each autoencoder of the task-agnostic images. It can automatically assign the testing data into MoE-Adapters or original CLIP to predict with either seen or unseen data.
@@ 3.3 Incremental Mixture-of-Experts Adapters
    Method
Aircraft [49]
Caltech101 [21]
CIFAR100 [38]
DTD [9]
EuroSAT [25]
Flowers [54]
Food [4]
MNIST [13]
OxfordPet [58]
Cars [37]
SUN397 [69]
Average
CLIP
   Zero-shot
24.3
88.4
68.2
44.6
54.9
71.0
88.5
59.4
89.0
64.7
65.2
65.3
   Full Fine-tune
62.0
95.1
89.6
79.5
98.9
97.5
92.7
99.6
94.7
89.6
81.8
89.2
   Fine-tune Adapter
56.8
92.6
89.4
79.0
98.4
97.0
92.9
99.2
94.1
89.1
82.7
88.3
Transfer
   Continual-FT
67.1
46.0
32.1
35.6
35.0
57.7
44.1
60.8
20.5
46.6
44.6
   LwF [42]
74.5
56.9
39.1
51.1
52.6
72.8
60.6
75.1
30.3
55.9
58.9
   iCaRL [61]
56.6
44.6
32.7
39.3
46.6
68.0
46.0
77.4
31.9
60.5
50.4
   LwF-VR [15]
77.1
61.0
40.5
45.3
54.4
74.6
47.9
76.7
36.3
58.6
57.2
   WiSE-FT [67]
73.5
55.6
35.6
41.5
47.0
68.3
53.9
69.3
26.8
51.9
52.3
   ZSCL [78]
86.0
67.4
45.4 50.4 69.1
87.6
61.8
86.8
60.1
66.8 68.1
@@
87.9 68.2
42.2
41.4
68.7
88.7
59.4
89.1 64.5
64.0
67.4 (-0.7)
   Ours
87.9 68.2 44.4
49.9
70.7 88.7
59.7
89.1 64.5 65.5 68.9 (+0.8)
Average
   Continual-FT
25.5
81.5
59.1
53.2
64.7
51.8
63.2
64.3
69.7
31.8
49.7
55.9
   LwF [42]
36.3
86.9
72.0
59.0
73.7
60.0
73.6
74.8
80.0
37.3
58.1
64.7
   iCaRL [61]
35.5
89.2
72.2
60.6
68.8
70.0
78.2
62.3
81.8
41.2
62.5
65.7
   LwF-VR [15]
29.6
87.7
74.4
59.5
72.4
63.6
77.0
66.7
81.2
43.7
60.7
65.1
   WiSE-FT [67]
26.7
86.5
64.3
57.1
65.7
58.7
71.1
70.5
75.8
36.9
54.6
60.7
   ZSCL [78]
45.1
92.0
80.1
64.3
79.5
81.6
89.6 75.2
88.9
64.7
68.0
75.4
@@
54.3
91.1
85.1 69.7
77.5
84.5 89.1
73.8
89.2 69.0
65.8
77.2(+1.8)
   Ours
50.2 91.9 83.1 69.4 78.9 84.0 89.1
73.7
89.3 67.7 66.9
76.7(+1.3)
Last
   Continual-FT
31.0
89.3
65.8
67.3
88.9
71.1
85.6
99.6
92.9
77.3
81.1
77.3
   LwF [42]
26.3
87.5
71.9
66.6
79.9
66.9
83.8
99.6
92.1
66.1
80.4
74.6
   iCaRL [61]
35.8
93.0
77.0
70.2
83.3
88.5
90.4
86.7
93.2
81.2
81.9
80.1
   LwF-VR [15]
20.5
89.8
72.3
67.6
85.5
73.8
85.7
99.6
93.1
73.3
80.9
76.6
   WiSE-FT [67]
27.2
90.8
68.0
68.9
86.9
74.0
87.6
99.6
92.6
77.8
81.3
77.7
   ZSCL [78]
40.6
92.2
81.3
70.5
94.8
90.5
91.9
98.7
93.9 85.3
80.2
83.6
@@
    Method
Aircraft [49]
Caltech101 [21]
CIFAR100 [38]
DTD [9]
EuroSAT [25]
Flowers [54]
Food [4]
MNIST [13]
OxfordPet [58]
Cars [37]
SUN397 [69]
Average
CLIP    Zero-shot
24.3
88.4
68.2
44.6
54.9
71.0
88.5
59.4
89.0
64.7
65.2
65.3
   5-shot Full Fine-tune
30.6
93.5
76.8
65.1
91.7
92.9
83.3
96.6
84.9
65.4
71.3
77.5
   5-shot Fine-tune Adapter
29.7
90.0
75.3
63.9
81.1
94.2
87.8
90.4
89.0
68.2
72.5
76.6
Transfer    Continual-FT
72.8
53.0
36.4
35.4
43.3
68.4
47.4
72.6
30.0
52.7
51.2
   LwF [42]
72.1
49.2
35.9
44.5
41.1
66.6
50.5
69.0
19.0
51.7
50.0
   LwF-VR [15]
82.2
62.5
40.1
40.1
56.3
80.0
60.9
77.6
40.5
60.8
60.1
   WiSE-FT [67]
77.6
60.0
41.3
39.4
53.0
76.6
58.1
75.5
37.3
58.2
57.7
@@
Average    Continual-FT
28.1
86.4
59.1
52.8
55.8
62.0
70.2
64.7
75.5
35.0
54.0
58.5
   LwF [42]
23.5
77.4
43.5
41.7
43.5
52.2
54.6
63.4
68.0
21.3
52.6
49.2
   LwF-VR [15]
24.9
89.1
64.2
53.4
54.3
70.8
79.2
66.5
79.2
44.1
61.6
62.5
   WiSE-FT [67] 32.0
87.7
61.0
55.8 68.1
69.3
76.8
71.5
77.6
42.0
59.3
63.7
   ZSCL [78]
28.2
88.6
66.5
53.5
56.3
73.4 83.1
56.4
82.4 57.5 62.9 64.4
@@
Last    Continual-FT
27.8
86.9
60.1
58.4
56.6
75.7
73.8
93.1
82.5
57.0
66.8
67.1
   LwF [42]
22.1
58.2
17.9
32.1
28.1
66.7
46.0
84.3
64.1
31.5
60.1
46.5
   LwF-VR [15]
22.9
89.8
59.3
57.1
57.6
79.2
78.3
77.7
83.6
60.1
69.8
66.9
   WiSE-FT [67] 30.8
88.9
59.6
60.3 80.9
81.7
77.1
94.9
83.2
62.8
70.0
71.9
   ZSCL [78]
26.8
88.5
63.7
55.7
60.2
82.1 82.6
58.6
85.9 66.7 70.4
67.4
   Ours 30.1 89.3 74.9 64.0 82.3 89.4 87.1
89.0
89.1 69.5 72.5 76.1(+4.2)
@@ 4.2 Comparison with State-of-the-art Methods
Method
Train Params $\downarrow$
GPU $\downarrow$
Times $\downarrow$
@@
ZSCL [78]
149.6M
26290MiB
3.94s/it
MoE-Adapters
  51.1M
19898MiB
1.37s/it
DDAS
  8.7M
  2461MiB
0.21s/it
Ours
  59.8M
22358MiB
1.58s/it
@@ E. Effectiveness of Router Selection in MoE-Adapters
Trans.
Avg.
Last
Trans.
Avg.
Last
Trans.
Avg.
Last
2
65.8
60.7
59.0
65.9
63.2
63.2
67.3
64.1
61.5
4
64.9
67.5
77.1
65.0
71.2
77.9
66.5
71.1
74.1
8
65.1 68.3 78.3
65.4
73.7
84.9
67.4
75.7
82.4
16
65.3
67.7
77.9
65.5
73.9
84.9
68.0 76.4 84.6
20
65.5
67.4
77.0
66.6 74.6 85.8
67.6
76.0
84.2
@@
Trans.
Avg.
Last
Trans.
Avg.
Last
ZSCL[78]
68.1
75.4
83.6
65.3
66.7
67.4
MAE
68.4
73.8
77.8
68.5
70.9
70.5
Smooth L1
68.3
76.9
84.9
69.0
72.9
72.6
MSE
68.9
76.7
85.0
68.9
76.3
76.1
@@
Aircraft [49]
Caltech101 [21]
CIFAR100 [38]
DTD [9]
EuroSAT [25]
Flowers [54]
Food [4]
MNIST [13]
OxfordPet [58]
Cars [37]
SUN397 [69]
Transfer
87.9
68.2
44.4
50.0
70.7
88.7
59.7
89.1
64.5
65.5
68.9
Aircraft 51.5
87.9
68.2
45.1
54.6
71.3
88.8
59.5
89.1
64.5
65.3
Caltech101
51.0
92.3
68.2
44.0
54.6
70.2
88.8
59.5
89.1
64.5
65.5
CIFAR100
50.0
91.5
86.7
44.2
44.4
70.8
88.8
59.8
89.1
64.5
65.5
DTD
50.4
92.0
86.5
78.6
45.9
70.7
88.8
59.8
89.1
64.5
65.5
EuroSAT
50.4
91.8
86.5
78.3
96.1
70.4
88.8
59.8
89.1
64.5
65.6
Flowers
50.3
92.3
86.3
79.1
95.7
95.9
88.7
59.8
89.1
64.5
65.6
Food
49.7
93.0
86.4
78.9
95.3
95.8
89.5
59.8
89.1
64.5
65.7
MNIST
49.7
92.7
86.3
79.0
95.5
95.6
89.5
98.3
89.1
64.5
65.6
OxfordPet
49.7
92.4
86.3
79.2
95.1
94.6
89.5
98.1
89.8
64.5
65.5
Cars
49.4
92.4
86.2
78.9
94.8
94.7
89.5
98.2
89.7
81.9
65.5
SUN397
49.8
92.2
86.1
78.1
95.7
94.3
89.5
98.1
89.9
81.6
80.0 85.0
Average
50.2
91.9
83.1
69.4
78.9
84.0
89.1
73.7
89.3
67.7
66.9
76.7
@@
Aircraft [49]
Caltech101 [21]
CIFAR100 [38]
DTD [9]
EuroSAT [25]
Flowers [54]
Food [4]
MNIST [13]
OxfordPet [58]
Cars [37]
SUN397 [69]
Transfer
87.9
68.2
42.4
41.4
68.7
88.7
59.4
89.1
64.5
64.0
67.4
Aircraft 54.3
87.9
68.2
45.1
54.6
71.3
88.8
59.5
89.1
64.5
65.3
Caltech101
54.2
92.0
68.2
40.7
54.6
67.7
88.7
59.5
89.1
64.5
63.5
CIFAR100
54.3
91.6
88.8
41.4
28.3
68.3
88.8
59.3
89.1
64.5
64.0
DTD
54.3
91.7
88.8
80.0
28.2
68.1
88.8
59.4
89.1
64.5
64.1
EuroSAT
54.3
91.7
88.8
79.9
98.0
68.1
88.8
59.4
89.1
64.5
64.0
Flowers
54.3
91.5
88.8
79.7
98.1
97.8
88.5
59.4
89.1
64.5
63.9
Food
54.3
91.0
88.8
80.0
98.1
97.8
89.7
59.3
89.1
64.5
63.8
MNIST
54.4
91.2
88.8
80.0
98.1
97.8
89.7
99.1
89.1
64.5
63.8
OxfordPet
54.3
90.8
88.8
79.8
98.1
97.6
89.6
99.1
89.6
64.5
63.4
Cars
54.2
90.8
88.8
80.2
98.1
97.5
89.6
99.1
89.5
89.2
63.8
SUN397
54.3
90.8
88.8
80.3
98.1
97.5
89.6
99.1
89.5
89.2
83.8 87.4
Average
54.3
91.0
85.1
69.7
77.5
84.5
89.1
73.8
89.2
69.0
65.8
77.2
@@
    Method
Cars [37]
Food [4]
MNIST [13]
OxfordPet [58]
Flowers [54]
SUN397 [69]
Aircraft [49]
Caltech101 [21]
DTD [9]
EuroSAT [25]
CIFAR100 [38]
Average
CLIP
   Zero-shot
64.7
88.5
59.4
89.0
71.0
65.2
24.3
88.4
44.6
54.9
68.2
65.3
   Full Fine-tune
89.6
92.7
99.6
94.7
97.5
81.8
62.0
95.1
79.5
98.9
89.6
89.2
   Fine-tune Adapter
89.1
92.9
99.2
94.1
97.0
82.7
56.8
92.6
79.0
98.4
89.4
88.3
Transfer
   Continual-FT
85.9
59.6
57.9
40.0
46.7
11.1
70.0
30.5
26.6
37.7
46.6
   LwF [42]
87.8
58.5
71.9
46.6
57.3
12.8
81.4
34.5
34.5
46.8
53.2
   iCaRL [61]
86.1
51.8
67.6
50.4
57.9
11.0
72.3
31.2
32.7
48.1
50.9
   LwF-VR [15]
88.2
57.0
71.4
50.0
58.0
13.0
82.0
34.4
29.3
47.6
53.1
   WiSE-FT [67]
87.2
57.6
67.0
45.0
54.0
12.9
78.6
35.5
28.4
44.3
51.1
   ZSCL [78]
88.3
57.5
84.7
68.1
64.8 21.1 88.2 45.3 55.2 68.2
64.1
@@
88.8 59.5 89.1 69.4 65.3
15.0
87.9 43.9 54.6 68.2
64.2(+0.1)
   Ours
88.8 59.5 89.1 69.9
64.4
18.1
86.9
43.7
54.6 68.2 64.3(+0.2)
Average
   Continual-FT
42.1
70.5
92.2
80.1
54.5
59.1
19.8
78.3
41.0
38.1
42.3
56.2
   LwF [42]
49.0
77.0
92.1
85.9
66.5
67.2
20.9
84.7
44.6
45.5
50.5
62.2
   iCaRL [61]
52.0
75.9
77.4
74.6
58.4
59.3
11.7
79.6
42.1
43.2
51.7
56.9
   LwF-VR [15]
44.9
75.8
91.8
85.3
63.5
67.6
16.9
84.9
44.0
40.6
51.3
60.6
   WiSE-FT [67]
52.6
79.3
91.9
83.9
63.4
65.2
23.3
83.7
45.4
40.0
48.2
61.5
   ZSCL [78]
81.7
91.3
91.1
91.0
82.9
72.5 33.6 89.7 53.3 62.8 69.9
74.5
@@
86.8
89.3
92.2
89.1
86.0 73.0
30.8
90.0
53.1
62.6 69.9 74.8(+0.3)
   Ours
84.9 89.9
89.3
91.4 86.2
72.2
33.4
89.4
53.3
61.4
69.9
74.7(+0.2)
Last
   Continual-FT
24.0
67.3
99.1
87.4
44.3
67.0
29.5
92.3
61.3
81.0
88.1
67.4
   LwF [42]
34.6
69.6
99.3
88.7
61.1
72.5
32.5
88.1
65.6
90.9
87.9
71.9
   iCaRL [61]
46.0
81.5
91.3
82.8
66.5
72.2
16.3
91.6
68.1
83.2
87.8
71.6
   LwF-VR [15]
27.4
61.2
99.4
86.3
60.6
70.7
23.4
88.0
61.3
84.3
88.1
68.2
   WiSE-FT [67]
35.6
76.9
99.5
89.1
62.1
71.8
27.8
90.8
67.0
85.6
87.6
72.2
   ZSCL [78]
78.2
91.1
97.6
92.5
87.4
78.2 45.0
92.3
72.7
96.2
86.3
83.4
@@
    Method
Cars [49]
Food [21]
MNIST [38]
OxfordPet [9]
Flowers [25]
SUN397 [54]
Aircraft [4]
Caltech101 [13]
DTD [58]
EuroSAT [37]
CIFAR100 [69]
Average
CLIP    Zero-shot
64.7
88.5
59.4
89.0
71.0
65.2
24.3
88.4
44.6
54.9
68.2
65.3
   5-shot Full Fine-tune
65.4
83.3
96.6
84.9
92.9
71.3
30.6
93.5
65.1
91.7
76.8
77.5
   5-shot Fine-tune Adapter
68.2
87.8
90.4
89.0
94.2
72.5
29.7
90.0
63.9
81.1
75.3
76.6
Transfer    Continual-FT
76.0
64.6
67.1
49.7
53.7
8.3
77.9
33.9
23.9
37.1
49.2
   LwF [42]
64.2
59.1
68.1
38.4
54.9
6.7
78.0
35.5
33.5
47.4
48.6
   LwF-VR [15]
80.1
55.4
77.7
50.4
61.4
9.1
83.5
40.1
31.5
54.8
54.4
   WiSE-FT [67]
77.3
60.0
76.9
54.2
58.0
11.1
81.8
37.6
31.7
48.1
53.7

### v2

@@ 2 Related Works
Figure 2: Overall framework of the proposed method. (a) At the training stage, CLIP’s image and text encoders $(\mathcal{F}_{I},\mathcal{F}_{T})$ take input samples from Task t. In each of transformer blocks, there is a MoE-Adapters, whose input is the tokens $\textbf{{x}}^{t}$ from MHSA. The router takes the task-specific [CLS] token $\textbf{{c}}^{t}$ as input and produces experts’ weights $W_{i}^{t}$ and $W_{j}^{t}$ to combine the expert’s output. DDAS is trained using only images via the MSE loss defined by Eq. 3. (b) At the inference stage, the proposed DDAS determines the data distribution by comparing the distribution $\{d^{t}\}_{t=1}^{T}$ in each autoencoder of the task-agnostic images. It can automatically assign the testing data into MoE-Adapters or original CLIP to predict with either seen or unseen data.
@@ 3.3 Incremental Mixture-of-Experts Adapters
    Method Aircraft [49] Caltech101 [21] CIFAR100 [38] DTD [9] EuroSAT [25] Flowers [54] Food [4] MNIST [13] OxfordPet [58] Cars [37] SUN397 [69] Average
CLIP    Zero-shot 24.3 88.4 68.2 44.6 54.9 71.0 88.5 59.4 89.0 64.7 65.2 65.3
   Full Fine-tune 62.0 95.1 89.6 79.5 98.9 97.5 92.7 99.6 94.7 89.6 81.8 89.2
   Fine-tune Adapter 56.8 92.6 89.4 79.0 98.4 97.0 92.9 99.2 94.1 89.1 82.7 88.3
Transfer    Continual-FT 67.1 46.0 32.1 35.6 35.0 57.7 44.1 60.8 20.5 46.6 44.6
   LwF [42] 74.5 56.9 39.1 51.1 52.6 72.8 60.6 75.1 30.3 55.9 58.9
   iCaRL [61] 56.6 44.6 32.7 39.3 46.6 68.0 46.0 77.4 31.9 60.5 50.4
   LwF-VR [15] 77.1 61.0 40.5 45.3 54.4 74.6 47.9 76.7 36.3 58.6 57.2
   WiSE-FT [67] 73.5 55.6 35.6 41.5 47.0 68.3 53.9 69.3 26.8 51.9 52.3
   ZSCL [78] 86.0 67.4 45.4 50.4 69.1 87.6 61.8 86.8 60.1 66.8 68.1
@@
   Ours 87.9 68.2 44.4 49.9 70.7 88.7 59.7 89.1 64.5 65.5 68.9 (+0.8)
Average    Continual-FT 25.5 81.5 59.1 53.2 64.7 51.8 63.2 64.3 69.7 31.8 49.7 55.9
   LwF [42] 36.3 86.9 72.0 59.0 73.7 60.0 73.6 74.8 80.0 37.3 58.1 64.7
   iCaRL [61] 35.5 89.2 72.2 60.6 68.8 70.0 78.2 62.3 81.8 41.2 62.5 65.7
   LwF-VR [15] 29.6 87.7 74.4 59.5 72.4 63.6 77.0 66.7 81.2 43.7 60.7 65.1
   WiSE-FT [67] 26.7 86.5 64.3 57.1 65.7 58.7 71.1 70.5 75.8 36.9 54.6 60.7
   ZSCL [78] 45.1 92.0 80.1 64.3 79.5 81.6 89.6 75.2 88.9 64.7 68.0 75.4
@@
   Ours 50.2 91.9 83.1 69.4 78.9 84.0 89.1 73.7 89.3 67.7 66.9 76.7(+1.3)
Last    Continual-FT 31.0 89.3 65.8 67.3 88.9 71.1 85.6 99.6 92.9 77.3 81.1 77.3
   LwF [42] 26.3 87.5 71.9 66.6 79.9 66.9 83.8 99.6 92.1 66.1 80.4 74.6
   iCaRL [61] 35.8 93.0 77.0 70.2 83.3 88.5 90.4 86.7 93.2 81.2 81.9 80.1
   LwF-VR [15] 20.5 89.8 72.3 67.6 85.5 73.8 85.7 99.6 93.1 73.3 80.9 76.6
   WiSE-FT [67] 27.2 90.8 68.0 68.9 86.9 74.0 87.6 99.6 92.6 77.8 81.3 77.7
   ZSCL [78] 40.6 92.2 81.3 70.5 94.8 90.5 91.9 98.7 93.9 85.3 80.2 83.6
@@
      Method Aircraft [49] Caltech101 [21] CIFAR100 [38] DTD [9] EuroSAT [25] Flowers [54] Food [4] MNIST [13] OxfordPet [58] Cars [37] SUN397 [69] Average
CLIP      Zero-shot 24.3 88.4 68.2 44.6 54.9 71.0 88.5 59.4 89.0 64.7 65.2 65.3
     5-shot Full Fine-tune 30.6 93.5 76.8 65.1 91.7 92.9 83.3 96.6 84.9 65.4 71.3 77.5
     5-shot Fine-tune Adapter 29.7 90.0 75.3 63.9 81.1 94.2 87.8 90.4 89.0 68.2 72.5 76.6
Transfer      Continual-FT 72.8 53.0 36.4 35.4 43.3 68.4 47.4 72.6 30.0 52.7 51.2
     LwF [42] 72.1 49.2 35.9 44.5 41.1 66.6 50.5 69.0 19.0 51.7 50.0
     LwF-VR [15] 82.2 62.5 40.1 40.1 56.3 80.0 60.9 77.6 40.5 60.8 60.1
     WiSE-FT [67] 77.6 60.0 41.3 39.4 53.0 76.6 58.1 75.5 37.3 58.2 57.7
@@
Average      Continual-FTSpace Space 28.1 86.4 59.1 52.8 55.8 62.0 70.2 64.7 75.5 35.0 54.0 58.5
     LwF [42] 23.5 77.4 43.5 41.7 43.5 52.2 54.6 63.4 68.0 21.3 52.6 49.2
     LwF-VR [15] 24.9 89.1 64.2 53.4 54.3 70.8 79.2 66.5 79.2 44.1 61.6 62.5
     WiSE-FT [67] 32.0 87.7 61.0 55.8 68.1 69.3 76.8 71.5 77.6 42.0 59.3 63.7
     ZSCL [78] 28.2 88.6 66.5 53.5 56.3 73.4 83.1 56.4 82.4 57.5 62.9 64.4
@@
Last      Continual-FT 27.8 86.9 60.1 58.4 56.6 75.7 73.8 93.1 82.5 57.0 66.8 67.1
     LwF [42] 22.1 58.2 17.9 32.1 28.1 66.7 46.0 84.3 64.1 31.5 60.1 46.5
     LwF-VR [15] 22.9 89.8 59.3 57.1 57.6 79.2 78.3 77.7 83.6 60.1 69.8 66.9
     WiSE-FT [67] 30.8 88.9 59.6 60.3 80.9 81.7 77.1 94.9 83.2 62.8 70.0 71.9
     ZSCL [78] 26.8 88.5 63.7 55.7 60.2 82.1 82.6 58.6 85.9 66.7 70.4 67.4
     Ours 30.1 89.3 74.9 64.0 82.3 89.4 87.1 89.0 89.1 69.5 72.5 76.1(+4.2)
@@ 4.2 Comparison with State-of-the-art Methods
Method Train Params $\downarrow$ GPU $\downarrow$ Times $\downarrow$
@@
ZSCL [78] 149.6M 26290MiB 3.94s/it
MoE-Adapters   51.1M 19898MiB 1.37s/it
DDAS   8.7M   2461MiB 0.21s/it
Ours   59.8M 22358MiB 1.58s/it
@@ E. Effectiveness of Router Selection in MoE-Adapters
Trans. Avg. Last Trans. Avg. Last Trans. Avg. Last
2 65.8 60.7 59.0 65.9 63.2 63.2 67.3 64.1 61.5
4 64.9 67.5 77.1 65.0 71.2 77.9 66.5 71.1 74.1
8 65.1 68.3 78.3 65.4 73.7 84.9 67.4 75.7 82.4
16 65.3 67.7 77.9 65.5 73.9 84.9 68.0 76.4 84.6
20 65.5 67.4 77.0 66.6 74.6 85.8 67.6 76.0 84.2
@@
Trans. Avg. Last Trans. Avg. Last
ZSCL[78] 68.1 75.4 83.6 65.3 66.7 67.4
MAE 68.4 73.8 77.8 68.5 70.9 70.5
Smooth L1 68.3 76.9 84.9 69.0 72.9 72.6
MSE 68.9 76.7 85.0 68.9 76.3 76.1
@@
Aircraft [49] Caltech101 [21] CIFAR100 [38] DTD [9] EuroSAT [25] Flowers [54] Food [4] MNIST [13] OxfordPet [58] Cars [37] SUN397 [69]
Transfer 87.9 68.2 44.4 50.0 70.7 88.7 59.7 89.1 64.5 65.5 68.9
Aircraft 51.5 87.9 68.2 45.1 54.6 71.3 88.8 59.5 89.1 64.5 65.3
Caltech101 51.0 92.3 68.2 44.0 54.6 70.2 88.8 59.5 89.1 64.5 65.5
CIFAR100 50.0 91.5 86.7 44.2 44.4 70.8 88.8 59.8 89.1 64.5 65.5
DTD 50.4 92.0 86.5 78.6 45.9 70.7 88.8 59.8 89.1 64.5 65.5
EuroSAT 50.4 91.8 86.5 78.3 96.1 70.4 88.8 59.8 89.1 64.5 65.6
Flowers 50.3 92.3 86.3 79.1 95.7 95.9 88.7 59.8 89.1 64.5 65.6
Food 49.7 93.0 86.4 78.9 95.3 95.8 89.5 59.8 89.1 64.5 65.7
MNIST 49.7 92.7 86.3 79.0 95.5 95.6 89.5 98.3 89.1 64.5 65.6
OxfordPet 49.7 92.4 86.3 79.2 95.1 94.6 89.5 98.1 89.8 64.5 65.5
Cars 49.4 92.4 86.2 78.9 94.8 94.7 89.5 98.2 89.7 81.9 65.5
SUN397 49.8 92.2 86.1 78.1 95.7 94.3 89.5 98.1 89.9 81.6 80.0 85.0
Average 50.2 91.9 83.1 69.4 78.9 84.0 89.1 73.7 89.3 67.7 66.9 76.7
@@
Aircraft [49] Caltech101 [21] CIFAR100 [38] DTD [9] EuroSAT [25] Flowers [54] Food [4] MNIST [13] OxfordPet [58] Cars [37] SUN397 [69]
Transfer 87.9 68.2 42.4 41.4 68.7 88.7 59.4 89.1 64.5 64.0 67.4
Aircraft 54.3 87.9 68.2 45.1 54.6 71.3 88.8 59.5 89.1 64.5 65.3
Caltech101 54.2 92.0 68.2 40.7 54.6 67.7 88.7 59.5 89.1 64.5 63.5
CIFAR100 54.3 91.6 88.8 41.4 28.3 68.3 88.8 59.3 89.1 64.5 64.0
DTD 54.3 91.7 88.8 80.0 28.2 68.1 88.8 59.4 89.1 64.5 64.1
EuroSAT 54.3 91.7 88.8 79.9 98.0 68.1 88.8 59.4 89.1 64.5 64.0
Flowers 54.3 91.5 88.8 79.7 98.1 97.8 88.5 59.4 89.1 64.5 63.9
Food 54.3 91.0 88.8 80.0 98.1 97.8 89.7 59.3 89.1 64.5 63.8
MNIST 54.4 91.2 88.8 80.0 98.1 97.8 89.7 99.1 89.1 64.5 63.8
OxfordPet 54.3 90.8 88.8 79.8 98.1 97.6 89.6 99.1 89.6 64.5 63.4
Cars 54.2 90.8 88.8 80.2 98.1 97.5 89.6 99.1 89.5 89.2 63.8
SUN397 54.3 90.8 88.8 80.3 98.1 97.5 89.6 99.1 89.5 89.2 83.8 87.4
Average 54.3 91.0 85.1 69.7 77.5 84.5 89.1 73.8 89.2 69.0 65.8 77.2
@@
    Method Cars [37] Food [4] MNIST [13] OxfordPet [58] Flowers [54] SUN397 [69] Aircraft [49] Caltech101 [21] DTD [9] EuroSAT [25] CIFAR100 [38] Average
CLIP    Zero-shot 64.7 88.5 59.4 89.0 71.0 65.2 24.3 88.4 44.6 54.9 68.2 65.3
   Full Fine-tune 89.6 92.7 99.6 94.7 97.5 81.8 62.0 95.1 79.5 98.9 89.6 89.2
   Fine-tune Adapter 89.1 92.9 99.2 94.1 97.0 82.7 56.8 92.6 79.0 98.4 89.4 88.3
Transfer    Continual-FT 85.9 59.6 57.9 40.0 46.7 11.1 70.0 30.5 26.6 37.7 46.6
   LwF [42] 87.8 58.5 71.9 46.6 57.3 12.8 81.4 34.5 34.5 46.8 53.2
   iCaRL [61] 86.1 51.8 67.6 50.4 57.9 11.0 72.3 31.2 32.7 48.1 50.9
   LwF-VR [15] 88.2 57.0 71.4 50.0 58.0 13.0 82.0 34.4 29.3 47.6 53.1
   WiSE-FT [67] 87.2 57.6 67.0 45.0 54.0 12.9 78.6 35.5 28.4 44.3 51.1
   ZSCL [78] 88.3 57.5 84.7 68.1 64.8 21.1 88.2 45.3 55.2 68.2 64.1
@@
   Ours 88.8 59.5 89.1 69.9 64.4 18.1 86.9 43.7 54.6 68.2 64.3(+0.2)
Average    Continual-FT 42.1 70.5 92.2 80.1 54.5 59.1 19.8 78.3 41.0 38.1 42.3 56.2
   LwF [42] 49.0 77.0 92.1 85.9 66.5 67.2 20.9 84.7 44.6 45.5 50.5 62.2
   iCaRL [61] 52.0 75.9 77.4 74.6 58.4 59.3 11.7 79.6 42.1 43.2 51.7 56.9
   LwF-VR [15] 44.9 75.8 91.8 85.3 63.5 67.6 16.9 84.9 44.0 40.6 51.3 60.6
   WiSE-FT [67] 52.6 79.3 91.9 83.9 63.4 65.2 23.3 83.7 45.4 40.0 48.2 61.5
   ZSCL [78] 81.7 91.3 91.1 91.0 82.9 72.5 33.6 89.7 53.3 62.8 69.9 74.5
@@
   Ours 84.9 89.9 89.3 91.4 86.2 72.2 33.4 89.4 53.3 61.4 69.9 74.7(+0.2)
Last    Continual-FT 24.0 67.3 99.1 87.4 44.3 67.0 29.5 92.3 61.3 81.0 88.1 67.4
   LwF [42] 34.6 69.6 99.3 88.7 61.1 72.5 32.5 88.1 65.6 90.9 87.9 71.9
   iCaRL [61] 46.0 81.5 91.3 82.8 66.5 72.2 16.3 91.6 68.1 83.2 87.8 71.6
   LwF-VR [15] 27.4 61.2 99.4 86.3 60.6 70.7 23.4 88.0 61.3 84.3 88.1 68.2
   WiSE-FT [67] 35.6 76.9 99.5 89.1 62.1 71.8 27.8 90.8 67.0 85.6 87.6 72.2
   ZSCL [78] 78.2 91.1 97.6 92.5 87.4 78.2 45.0 92.3 72.7 96.2 86.3 83.4
@@
      Method Cars [49] Food [21] MNIST [38] OxfordPet [9] Flowers [25] SUN397 [54] Aircraft [4] Caltech101 [13] DTD [58] EuroSAT [37] CIFAR100 [69] Average
CLIP      Zero-shot 64.7 88.5 59.4 89.0 71.0 65.2 24.3 88.4 44.6 54.9 68.2 65.3
     5-shot Full Fine-tune 65.4 83.3 96.6 84.9 92.9 71.3 30.6 93.5 65.1 91.7 76.8 77.5
     5-shot Fine-tune Adapter 68.2 87.8 90.4 89.0 94.2 72.5 29.7 90.0 63.9 81.1 75.3 76.6
Transfer      Continual-FT 76.0 64.6 67.1 49.7 53.7 8.3 77.9 33.9 23.9 37.1 49.2
     LwF [42] 64.2 59.1 68.1 38.4 54.9 6.7 78.0 35.5 33.5 47.4 48.6
     LwF-VR [15] 80.1 55.4 77.7 50.4 61.4 9.1 83.5 40.1 31.5 54.8 54.4
     WiSE-FT [67] 77.3 60.0 76.9 54.2 58.0 11.1 81.8 37.6 31.7 48.1 53.7